OpenStudy (anonymous):

If G(x) = 3x + 1, then G -1(1) is answer choices 4, 1/3 ,0

myininaya (myininaya):

So first, if you want to find \[G^{-1}(1)=? \] Then all you real need to do is find for what value ? that G(?)=1. That is one way to do it. You could actually start by finding the inverse of G. So do you know how to solve the following equation for ? . G(?)=1

myininaya (myininaya):

Find G(?) first. Just take the G(x)=3x+1 and replace all the x's with ?'s

myininaya (myininaya):

Then solve G(?)=1 for ?

OpenStudy (anonymous):

This is the question

myininaya (myininaya):

right. So you know how to find G(?) ?

OpenStudy (anonymous):

no

myininaya (myininaya):

Do you see the G(x)=3x+1 thing? To find G(?) replace all the x's in that above function with ?'s

OpenStudy (anonymous):

alright G(?)=3?+1

myininaya (myininaya):

now solve G(?)=1 for ?

myininaya (myininaya):

replace G(?) with 3*?+1 like so G(?)=1 3*?+1=1

OpenStudy (anonymous):

its 1/3 right

myininaya (myininaya):

solve that equation for ?

OpenStudy (anonymous):

0

myininaya (myininaya):

right \[\text{since } G(0)=1 \text{ then } G^{-1}(1)=0\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Given that F(x)=x 2+2, evaluate F(2). 0 2 6

myininaya (myininaya):

F(2) means to replace the x in F(x)=x 2 +2 with 2.

OpenStudy (anonymous):

sorry wrong one

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

replace the x with what it says to replace it with

myininaya (myininaya):

then evaluate

OpenStudy (anonymous):

its 6

OpenStudy (anonymous):

2^+2=6

myininaya (myininaya):

Just like when I told you to evaluate G(?) given G(x)=3x+1 you did G(?)=3*?+1 so if F(x)=x^2+2 then F(2)=2^2+2 right

myininaya (myininaya):

well do you see any multiplication or addition?

OpenStudy (anonymous):

no

myininaya (myininaya):

so that means you can eliminate what choices?

OpenStudy (anonymous):

both of those

myininaya (myininaya):

what choices are left?

OpenStudy (anonymous):

distributive and sym

myininaya (myininaya):

do you know the distributive property?

OpenStudy (anonymous):

can you do me a favor

myininaya (myininaya):

What is the favor?

OpenStudy (anonymous):

can you sign into this site under my name and do this for me ill pay or you could be kind hearted and finish this moduel for me your a math wiz

myininaya (myininaya):

That is cheating though. I cannot do your work. It is not helping you. It would only be damaging you for future tests, quizzes, and other homework. Please look at the CoC http://openstudy.com/code-of-conduct and the terms and conditions http://openstudy.com/terms-and-conditions . These are both things you agreed to when you signed up on OpenStudy.

OpenStudy (anonymous):

youre right i need to do this but damn seems like nobody wants to help anybody on here

myininaya (myininaya):

I was trying to help you...

OpenStudy (anonymous):

yeah but youre gonna leave sonner or later

OpenStudy (anonymous):

wrong one

myininaya (myininaya):

Here is a hint: \[x^0 =1 \text{ note: assuming } x \neq 0\]

OpenStudy (anonymous):

its 1

myininaya (myininaya):

yep anything with the power 0 is 1 except for when you have 0^0

myininaya (myininaya):

you are suppose to go under the assumption a+b is not 0 so (a+b)^0 would be 1

myininaya (myininaya):

hint: \[\frac{1}{x^n}=x^{-n}\]

myininaya (myininaya):

\[\frac{1}{(something)^n}=(something)^{-n}\]

OpenStudy (anonymous):

1/6?

myininaya (myininaya):

I'm seeing this problem: \[\frac{1}{(xy)^1}\]

myininaya (myininaya):

Are you doing a different one...

OpenStudy (anonymous):

yes thats the right problem

myininaya (myininaya):

look at what I wrote above what is the something and what is the n value?

myininaya (myininaya):

compare 1/(xy)^1 to 1/(something)^n

myininaya (myininaya):

determine what something is and what the n is and then replace it on the side of that equation i wrote

myininaya (myininaya):

\[\frac{1}{(something)^n}=(something)^{-n}\]

myininaya (myininaya):

\[\frac{1}{(xy)^1}=( ? )^{- *}\] your job is to replace that ? and to replace the *

OpenStudy (anonymous):

1x and i dont know the last part is it -2

myininaya (myininaya):

I see no 2's in 1/(xy)^1

myininaya (myininaya):

what happen to the y

OpenStudy (anonymous):

yeah youre right

OpenStudy (anonymous):

wym

myininaya (myininaya):

\[\frac{1}{(xy)^n}=\frac{1}{(something)^n}\] Compare both sides of this equation What would the something have to be and what would the n have to be for both of the sides to be the same as each other. what is in the something's place what is in the n's place

OpenStudy (anonymous):

1

myininaya (myininaya):